Hopp Company requires a minimum cash balance of $3,100. When the company expects a cash​ deficiency, it borrows the exact amount required on the first of the month. Expected excess cash is used to repay any amounts owed. Interest owed from the previous​ month's principal balance is paid on the first of the month at 17​% per year. The company has already completed the budgeting process for the first quarter for cash receipts and cash payments for all expenses except interest.

Begin by preparing the cash budget for January then prepare the cash budget for February and March. Finally, prepare the totals for the quarter.

Hip Company

Cash Budget

For the Three Months Ended March 31

January

February

March

Total

Beginning cash balance

$3,100

Cash receipts

22,500

29,500

42,500

94,500

Cash available

25,600

Cash payments:

All expenses except interest

31,000

34,000

36,000

101,000

Interest expense

0

Total cash payments

31,000

Ending cash balance before financing

Minimum cash balance desired

(3,100)

Projected cash excess (deficiency)

Financing:

Borrowing

Principal repayments

Total effects of financing

Ending cash balance
Answer #1
Interest expense in February $1,445 (17%*8500)
Interest expense in March $2,456 (17%*(8500+5945)
January February March Total
a Beginning Cash Balance $3,100 $3,100 $3,100
b Cash Receipts $22,500 $29,500 $42,500 $94,500
c=a+b Total Cash Available $25,600 $32,600 $45,600
d All expenses except interest $31,000 $34,000 $36,000 $101,000
e Interest expense $0 $1,445 $2,456 $3,901
f=d+e Total expense $31,000 $35,445 $38,456 $104,901
g=c-f Excess/(deficit ) of Cash ($5,400) ($2,845) $7,144
h Minimum balance required $3,100 $3,100 $3,100
i Borrowing $8,500 $5,945 $0 $14,445
j Repayment $0 $0 ($4,044) ($4,044)
k=i+j Total effect of financing $8,500 $5,945 ($4,044) $10,401
l=g+k Ending Cash Balance $3,100 $3,100 $3,100